# Monotonic Path-Specific Effects: Application to Estimating Educational Returns

### v1

@@ 4 Empirical Analysis
5 Conclusion
Although my empirical motivation is the estimation of educational returns, the proposed framework applies widely to a range of demographic and organizational settings characterized by “state dependency” between treatment and mediators. This characteristic is particularly salient in demographic phenomena, which often involve sequential transitions over the life course. Certain demographic events are rigid in their monotonicity as a result of their definition. For example, researchers may be interested in discerning the degree to which positive effects of marriage on outcomes such as earnings and life satisfaction are undermined by the negative effects of divorce and separation (and, in turn, their mitigation via re-marriage) (Kenney, 2004; Sweeney and Phillips, 2004). Divorce can be “attained” only by individuals who are already married. Similarly, the effect of parenthood on earnings can be seen as operating directly, through the effect of having a first child net of subsequent children, as well as operating indirectly through the effects of having multiple children, transitions which are clearly monotonic in nature. A similar perspective may be taken in a criminal justice context: the total effect of early-stage police contact (such as being searched for contraband) on educational and socio-psychological outcomes can be decomposed into path-specific effects via subsequent arrest and incarceration (Weaver and Lerman, 2010; Kirk and Sampson, 2013; Sugie and Turney, 2017).
Finally, although in this paper I have considered a decomposition of the average treatment effect for the case of binary monotonic mediators, as shown in Supplementary Material G, the framework could straightforwardly be extended to accommodate categorical transitions. Given the heterogeneity of higher-educational trajectories in the US, such an extension would prove useful for modeling the relative payoffs to distinct educational pathways.
SUPPLEMENTARY MATERIAL
Supplementary Materials (To Appear Online):
This document (PDF) contains: (A) details about a parametric, regression-with-residuals estimation procedure; (B) a simulation study, (C) a sensitivity analysis, (D) further information on sample construction and the data used, (E-F) results with alternative sample restrictions and definitions of earnings; (G) a discussion of extending the decomposition to multivalued, discrete mediators; (H) a description of the EIFs used in the empirical illustration, and (I) proofs and technical details.
@@ 5 Conclusion
1 Parametric, regression-with-residuals (RWR) estimation
@@
      compute least squares estimates of equations 11 and 12, using estimates of $X^{\perp}$ and $Z_{k}^{\perp}$.
@@
2 A simulation study
In this section, I evaluate the finite-sample performance of my two estimation procedures via a simulation experiment. Specifically, I compare how the DML estimator proposed in Section 3 (as well as the parametric, RWR estimator described in Appendix 1) perform under different degrees of misspecification. Without loss of generality, I focus on the single-mediator setting, and focus on the path-specific effect $A\to M\to Y$. I generate simulations of observed data $O=(X_{1},X_{2},A,Z,M,Y)$ as follows:
@@ 2 A simulation study
$A$ $\sim\text{Bern}\big{(}\text{\text{logit}}^{-1}\big{[}(1,X)\beta_{A}\big{]}\big{)}$
$Z|A=1$ $\sim\text{Bern}\big{(}\text{\text{logit}}^{-1}\big{[}(1,X)\beta_{Z}\big{]}\big{)}$
$M|A=1\sim\text{Bern}\big{(}\text{\text{logit}}^{-1}\big{[}(1,X,Z)\beta_{M}\big{]}\big{)}$ $\quad M|A=0=0$
$Y$ $\sim\text{N}\big{(}(1,A,X,AZ,AM)\beta_{Y},1\big{)}.$
The coefficients $(\beta_{X_{1}},\beta_{X_{2}},\beta_{Y})$ are drawn from a $\text{Unif}\big{(}-1,1\big{)}$ , while the coefficient $\beta_{A}$ is drawn from a $\text{Unif}\big{(}-0.5,0.5\big{)}$ distribution. In order to test how the DML and RWR methods perform when the relevant models are misspecified, I also construct transformations of the observed covariates ($X_{i}^{*},Z_{i}^{*}$) as follows, employing a similar setup to Kang and Schafer [2007]:
@@
[…]
For each simulated dataset, I construct two estimates of the path-specific effect $\theta_{1}$ ($A\to M\to Y$) by estimating the parameter set ($\tau_{0}$, $\Delta_{0},\Delta_{1},\pi_{1})$ via the RWR and DML procedures described in Section 3. Standard errors for the coverage rates are computed via the estimated variance of the estimated EIFs for the DML approach, and via the nonparametric bootstrap with 1000 replications for the RWR procedure. For the DML estimator, for each component involved in the decomposition, I construct a Neyman-orthogonal “signal” using its EIF. The recentered EIFs for each component are shown below:
$M^{*}(1)$ $=\gamma_{1}(X)+\frac{\mathbb{I}(A=1)}{\pi_{0}(X,1)}(M-\gamma_{1}(X)),$
$Y^{*}(1,m_{1})$ $=\nu_{1}(X,m)+\frac{\mathbb{I}(A=1)\mathbb{I}(M=m)}{\pi_{0}(X,1)\pi_{1}(X,m_{1})}(Y-\mu_{1}(X,Z,m))$
$+\frac{\mathbb{I}(A=1)}{\pi_{0}(X,1)}(\mu_{1}(X,Z,m)-\nu_{1}(X,m)),\;\text{for }m\in\{0,1\}$
@@
$\pi_{0}(X,a)$ $\triangleq\mathrm{Pr}[A=a\mid X]$
$\pi_{1}(X,m_{1})$ $\triangleq\mathrm{Pr}[M=m\mid X,A=1]$
$\gamma_{1}(X)$ $\triangleq\mathbb{E}[M\mid X,A=1]$
$\mu_{0}(X,a)$ $\triangleq\mathbb{E}[Y|X,A=a]$
$\mu_{1}(X,Z,m)$ $\triangleq\mathbb{E}[Y|X,A=1,Z,M=m]$
$\nu_{1}(X,m)$ $\triangleq\mathbb{E}[\mu_{1}(X,Z,m)|X,A=1].$
I run 1000 replications of this DGP and compute the average bias, the root mean square error (RMSE), and the coverage of nominal 95% confidence intervals for sample sizes of 250, 500, and 1000 and using either the ”correctly specified” covariates $\left(X_{1},X_{2},Z\right)$ and the ”incorrectly specified”, transformed versions $\left(X_{1}^{*},X_{2}^{*},Z^{*}\right)$. I calculate the true value of $\theta_{1}$ by recovering the true values of the parameter set ($\tau_{0}$, $\Delta_{0},\Delta_{1},\pi_{1})$ in each Monte Carlo simulation.
@@
Figure 3 presents the results from this simulation exercise. Under correctly specified models, the DML and RWR estimators perform similarly, with each displaying low bias, low RMSE and close to nominal coverage at all sample sizes. Under incorrectly specified models, however, the DML and RWR approaches diverge in performance. Whereas the DML estimator under an incorrect feature space performs similarly to when it is used on the correct feature space, the RWR estimator performs much more poorly, displaying a large amount of bias that in fact grows with the sample size, a large RMSE, and coverage rates that are not close to nominal. In short, when the models are correctly specified, both the parametric and semiparametric approaches perform well; the strong performance of a semiparametric approach compared with a parametric estimation strategy becomes clearer under model misspecification.
@@
Figure 3: Bias, RMSE, and coverage of DML and RWR estimators for $n=250,500,1000$. The left panel shows the performance of the DML and RWR estimators when the correct feature matrix is supplied to the estimators; the right panel shows the performance of the two estimators when an incorrect feature matrix is supplied to the estimators, as described in the main text.
3 Sensitivity analysis
@@
4 Further details on variable construction and education groups
@@
Note: The category “High School Non-Completers” captures all individuals who attended high school but did not obtain a high school diploma; “High School Graduates” refers to those individuals who graduated high school, regardless of their subsequent educational experiences (i.e., whether or not they proceeded to college); “College Goers” refers to individuals who attended a 4-year college, irrespective of whether they completed their degree; “BA Completers” denotes individuals who completed a Bachelor’s degree, while “Grad. School Goers” captures individuals who participated in a graduate-level degree program. A small constant of \$1,000\\$1,000 is added to observed earnings before taking the log. All statistics are computed with a monotonicity assumption imposed on the observed data (i.e. such that all individuals who complete a given educational level are coded as having completed all prior levels). All statistics are calculated using NLSY97 sampling weights, and standard errors are in parentheses.
5 Results without imputation of missing covariates

### v2

@@ 6 Tables and figures
Figure 1: Nested decomposition of the Average Treatment Effect (ATE) into direct and sequential indirect pathways: general hierarchical decomposition (Panel A) and proposed monotonic decomposition (Panel B).
@@
A Parametric, regression-with-residuals (RWR) estimation
@@
      compute least squares estimates of equations 15 and 16, using estimates of $X^{\perp}$ and $Z_{k}^{\perp}$.
@@
B A simulation study
In this section, I evaluate the finite-sample performance of my two estimation procedures via a simulation experiment. Specifically, I compare how the DML estimator proposed in Section 3 (as well as the parametric, RWR estimator described in Appendix A) perform under different degrees of misspecification. Specifically, I consider the setting of two causally ordered monotonic mediators, with post-treatment confounding. I generate simulations of observed data $O=(X_{1},X_{2},A,Z,M,Y)$ as follows:
@@ B A simulation study
$A$ $\sim\text{Bern}\big(g^{-1}\big[(1,X)\beta_{A}\big]\big)$
$Z|A=1$ $\sim\text{N}\big[(1,X)\beta_{Z}\big]\big)$
$M_{1}|A=1\sim\text{Bern}\big(g^{-1}\big[(1,X,Z)\beta_{M_{1}}\big]\big)$ $\quad M|A=0=0$
$M_{2}|M_{1}=1\sim\text{Bern}\big(g^{-1}\big[(1,X,Z)\beta_{M_{2}}\big]\big)$ $\quad M_{2}|M_{1}=0=0$
$Y$ $\sim\text{N}\big((1,A,X,AZ,AM_{1},AM_{1}M_{2})\beta_{Y},1\big).$
The coefficients $(\beta_{X_{1}},\beta_{X_{2}},\beta_{Y})$ are drawn from a $\text{Unif}\big(-1,1\big)$ distribution, while the coefficient $\beta_{A}$ is drawn from a $\text{Unif}\big(-0.5,0.5\big)$ distribution. Further, $X=(X_{1},X_{2})$. In order to test how the DML and RWR methods perform when the relevant models are misspecified, I also construct transformations of the observed covariates ($X^{*}$) as follows, employing a similar setup to Kang and Schafer [2007]:
@@
When evaluating the DML estimation procedure, I set $g$ to be the logistic link: $g^{-1}(x)\;=\;\frac{\exp(x)}{1+\exp(x)}$. While the DML estimator is agnostic to the functional form of $g$: all nuisance functions—including propensity scores and mediator models—are estimated using flexible machine-learning methods, the RWR estimator relies on parametric linear regressions for both mediator and outcome models, when evaluating the RWR estimator I set $g$ to be the identity link.  [^7] More subtly, even if the outcome model is linear in $(M_{1},M_{2})$ by construction, the regression of $Y$ on $(X,A,M_{1},Z)$ used by RWR is generally misspecified when mediator models are nonlinear. To see this, note that $\mathbb{E}[Y\mid X,A=1,M_{1}=m,Z]=\mathbb{E}\!\left[\mathbb{E}[Y\mid X,A=1,M_{1}=m,M_{2},Z]\;\middle|\;X,A=1,M_{1}=m,Z\right],$ which involves integrating a linear function of $M_{2}$ with respect to a nonlinear conditional distribution of $M_{2}\mid X,Z,A=1,M_{1}=1$. Unless $\mathbb{E}[M_{2}\mid X,Z,A=1,M_{1}=1]$ is linear in $X$, this marginal conditional expectation is itself nonlinear in $X$. Thus, even when the structural outcome model is linear, the reduced-form outcome regression used by RWR will misspecified unless $g$ is the identity link.
For each simulated dataset, I construct two estimates of the path-specific effects ($\theta_{0},\theta_{1},\theta_{2}$) via the RWR and DML procedures described in Section 3. Standard errors for the coverage rates are computed via the estimated variance of the estimated EIFs for the DML approach, and via the nonparametric bootstrap with 250 replications for the RWR procedure. For the DML estimator, for each component involved in the decomposition, I construct a Neyman-orthogonal “signal” using its EIF. The recentered EIFs for each component are shown below:
$M_{1}^{*}(1)$ $=\gamma_{1}(X)+\frac{\mathbb{I}(A=1)}{\pi_{0}(X,1)}\bigl(M_{1}-\gamma_{1}(X)\bigr),$
$M_{2}^{*}(1,1)$ $=\gamma_{2}(X)+\frac{\mathbb{I}(A=1)\mathbb{I}(M_{1}=1)}{\pi_{0}(X,1)\,\pi_{1}(X,Z,1)}\bigl(M_{2}-\gamma_{2}(X)\bigr),$
$Y^{*}(a)$ $=\mu_{0}(X,a)+\frac{\mathbb{I}(A=a)}{\pi_{0}(X,a)}\bigl(Y-\mu_{0}(X,a)\bigr),\qquad a\in\{0,1\},$
$Y^{*}(1,m_{1})$ $=\nu_{1}(X,m_{1})+\frac{\mathbb{I}(A=1)\mathbb{I}(M_{1}=m_{1})}{\pi_{0}(X,1)\,\pi_{1}(X,Z,m_{1})}\bigl(Y-\mu_{1}(X,Z,m_{1})\bigr)$
$\qquad+\frac{\mathbb{I}(A=1)}{\pi_{0}(X,1)}\bigl(\mu_{1}(X,Z,m_{1})-\nu_{1}(X,m_{1})\bigr),\qquad m_{1}\in\{0,1\},$
$Y^{*}(1,1,m_{2})$ $=\nu_{2}(X,m_{2})+\frac{\mathbb{I}(A=1)\mathbb{I}(M_{1}=1)\mathbb{I}(M_{2}=m_{2})}{\pi_{0}(X,1)\,\pi_{1}(X,Z,1)\,\pi_{2}(X,Z,m_{2})}\bigl(Y-\mu_{2}(X,Z,m_{2})\bigr)$
$\qquad+\frac{\mathbb{I}(A=1)\mathbb{I}(M_{1}=1)}{\pi_{0}(X,1)\,\pi_{1}(X,Z,1)}\bigl(\mu_{2}(X,Z,m_{2})-\nu_{2}(X,m_{2})\bigr),\qquad m_{2}\in\{0,1\}.$
@@
$\pi_{0}(X,a)$ $\triangleq\Pr(A=a\mid X),$
$\pi_{1}(X,Z,m_{1})$ $\triangleq\Pr(M_{1}=m_{1}\mid X,Z,A=1),$
$\pi_{2}(X,Z,m_{2})$ $\triangleq\Pr(M_{2}=m_{2}\mid X,Z,A=1,M_{1}=1),$
$\gamma_{1}(X)$ $\triangleq\mathbb{E}[M_{1}\mid X,A=1],$
$\gamma_{2}(X)$ $\triangleq\mathbb{E}[M_{2}\mid X,A=1,M_{1}=1],$
$\mu_{0}(X,a)$ $\triangleq\mathbb{E}[Y\mid X,A=a],$
$\mu_{1}(X,Z,m_{1})$ $\triangleq\mathbb{E}[Y\mid X,A=1,Z,M_{1}=m_{1}],$
$\nu_{1}(X,m_{1})$ $\triangleq\mathbb{E}\!\left[\mu_{1}(X,Z,m_{1})\mid X,A=1\right],$
$\mu_{2}(X,Z,m_{2})$ $\triangleq\mathbb{E}[Y\mid X,A=1,Z,M_{1}=1,M_{2}=m_{2}],$
$\nu_{2}(X,m_{2})$ $\triangleq\mathbb{E}\!\left[\mu_{2}(X,Z,m_{2})\mid X,A=1,M_{1}=1\right].$
I run 1000 replications of this DGP and compute the average bias and coverage of nominal 95% confidence intervals for sample sizes of 1000, 1500 and 2000 and using either the "correctly specified" covariates $\left(X_{1},X_{2},Z\right)$ and the "incorrectly specified", transformed versions $\left(X_{1}^{*},X_{2}^{*},Z\right)$. I calculate the true value of $\theta_{1}$ by recovering the true values of the parameter set ($\theta_{0}$, $\theta_{1},\theta_{2}$) in each Monte Carlo simulation. When both the outcome model and the mediator models are linear probability models, these quantities can be derived analytilally. For example, under linearity, the continuation effect associated with the second mediator coincides with the coefficient on $M_{2}$ in the outcome model, and used to calculate the MPSE via $M_{2}$. However, under a nonlinear DGP, such closed-form expressions for the $\theta_{k}$ terms no longer exist because the mediator counterfactuals are nonlinear functions of $(X,Z)$. In this case, I recover the true values of $\left(\theta_{0},\theta_{1},\theta_{2}\right)$ by Monte Carlo integration under the known data-generating process. Specifically, for each replication, I simulate a large population from the structural equations under the relevant interventions in order to recover the counterfactual mediator probabilities. I then plug these counterfactual probabilities into the corresponding linear expressions implied by the outcome model to obtain the true values of the path-specific effects. These Monte Carlo quantities are treated as the ground truth against which finite-sample bias and coverage are evaluated.
Figure 4 presents the results from this simulation exercise. Under correctly specified models, the DML and RWR estimators perform similarly, with each displaying low bias and close to nominal coverage at all sample sizes. In particular, both estimators exhibit negligible finite-sample bias for $\theta_{0}$, $\theta_{1}$, and $\theta_{2}$, with absolute biases on the order of $10^{-3}$ or smaller, and coverage rates close to the nominal $95\%$ level. Increasing the sample size from $n=1000$ to $n=2000$ leads to only modest improvements.
@@
Under incorrectly specified models, however, the performance of the two estimators diverges sharply. The DML estimator remains stable: even when supplied with a misspecified feature space, absolute biases remain small for all MPSEs, and coverage rates remain close to nominal, typically between $92\%$ and $95\%$. By contrast, the RWR estimator performs much more poorly, displaying a large amount of bias that in fact grows with the sample size, a large RMSE, and coverage rates that are not close to nominal. In short, when the models are correctly specified, both the parametric and semiparametric approaches perform well; the strong performance of a semiparametric approach compared with a parametric estimation strategy becomes clearer under model misspecification. By contrast, the RWR estimator performs poorly under model misspecification. These failures become more pronounced as the sample size grows. For the direct effect $\theta_{0}$, bias increases dramatically with $n$, rising from approximately $0.32$ at $n=1000$ to over $1.1$ at $n=2000$. These biases are accompanied by severe inferential distortions: while coverage for $\theta_{0}$ remains artificially high due to the estimator centering far from the truth, coverage for $\theta_{2}$ declines substantially, reaching approximately $75\%$ at $n=2000$.
In sum, when the relevant models are correctly specified, both the parametric and semiparametric approaches perform well. Under misspecification, however, the advantages of the semiparametric estimation strategy become clear.
It is worthwhile briefly clarifying how the proposed monotonic path-specific effect (MPSE) decomposition relates to—and departs from—conventional mediation analyses of the ATE with multiple causally ordered mediators. Without mediator monotonicity, the ATE admits an algebraic decomposition into a collection of path-specific effects (PSEs) corresponding to the causal paths $A\to Y$ and $A\to M_{k}\rightsquigarrow Y$ for $k=1,\dots,K$ [Avin et al., 2005, Daniel et al., 2015, Zhou, 2022b]. However, when mediators are causally ordered, only composite effects of the form $A\to M_{k}\rightsquigarrow Y$—which aggregate all downstream pathways from $M_{k}$ to YY—are generally identifiable.
Under mediator monotonicity, this structure collapses in an important way. Because later mediators are deterministically zero whenever earlier mediators are not realized, path-specific effects of the form $A\to M_{k}\to Y$ for $k\geq 2$ are identically zero (see Section 2.3 in the main text). Consequently, the general PSE decomposition simplifies to a single-mediator natural effect decomposition in which the total indirect effect operates exclusively through the full causal chain $A\to M_{1}\to\cdots\to M_{k}\to Y$. This collapse has two implications for the simulation results. First, it explains why conventional multi-mediator decompositions do not provide a meaningful benchmark in this setting: once monotonicity is imposed, the distinction between multiple indirect paths disappears, and the estimand reduces to a single composite mediation effect. Second, and more importantly, even this reduced decomposition is not identifiable under standard mediation assumptions when intermediate confounders are present. In the simulation design considered here, these assumptions are deliberately violated by allowing for observed post-treatment confounders $Z$ that affect both later mediators and the outcome.
@@
Figure 4: Bias, RMSE, and coverage of DML and RWR estimators for $n=1000,1500,2000$. The red dots show the performance of the DML and RWR estimators when the correct feature matrix is supplied to the estimators; the blue dots show the performance of the two estimators when an incorrect feature matrix is supplied to the estimators.
C Sensitivity analysis
@@
Another important assumption underpinning the proposed decomposition is positivity. Since positivity is a key requirement for identifying each transition featuring in the MPSE decomposition (Assumption 5), it is important to assess the plausibility of this assumption. Positivity concerns are most plausible at the graduate-school transition. By the time individuals reach this stage, they have already passed through several earlier educational transitions. For instance, marginal BA completers or respondents with weaker academic records may have almost no chance of enrolling in graduate school, which raises concerns about overlap at this final stage.
Figure 6 displays the empirical distribution of the estimated propensity score for graduate-school attendance among the BA-completer risk set. The distribution shows substantial dispersion with no mass near 0 or 1. The minimum estimated probability is 0.11, and the 99th percentile is 0.68. Thus, the estimated propensities are bounded away from zero across the covariate support, and we do not observe a subgroup with near-zero model-implied probability of graduate education. This suggests no major violations of the positivity assumption at this final and most restrictive stage of the educational sequence.
@@
Figure 6: Distribution of the estimated graduate-school propensity score $\Pr(M_{3}=1\mid X,Z,A=1,M_{1}=1,M_{2}=1)$ among BA completers. The distribution shows substantial overlap without concentration near 0 or 1, supporting the plausibility of the positivity assumption for the graduate-school transition.
D Further details on variable construction and education groups
@@
Note: The category "High School Non-Completers" captures all individuals who attended high school but did not obtain a high school diploma; "High School Graduates" refers to those individuals who graduated high school, regardless of their subsequent educational experiences (i.e., whether or not they proceeded to college); "College Goers" refers to individuals who attended a 4-year college, irrespective of whether they completed their degree; "BA Completers" denotes individuals who completed a Bachelor’s degree, while "Grad. School Goers" captures individuals who participated in a graduate-level degree program. A small constant of $\mathdollar 1,000$ is added to observed earnings before taking the log. All statistics are computed with a monotonicity assumption imposed on the observed data (i.e. such that all individuals who complete a given educational level are coded as having completed all prior levels). All statistics are calculated using NLSY97 sampling weights, and standard errors are in parentheses.
E Results without imputation of missing covariates
In the main text, I report estimates of the MPSE decomposition for the ATE of high school attendance on logged annual earnings for the full sample of NLSY97 respondents with non-missing educational information and non-missing earnings ($N=7,305$). A very large number (approximately $50\%$) of these respondents are missing information on one or more of the covariates $(X,Z)$ used in the models in order to identify the decomposition components. Table 4 summarizes missingness patterns for the background covariates $X$ by respondents’ educational attainment - which are all self-reported by respondents. Missingness is generally modest for most pretreatment covariates, though certain variables (e.g., household net worth and ASVAB percentile) exhibit higher rates of non-response (at $25\%$ and $20\%$ of the sample overall, respectively). Non-response is more common among respondents with lower levels of schooling.
Table 5 clarifies the nature of missingness in the intermediate covariates $Z$ (STEM major and college GPA). These intermediate covariates are observed only for respondents who attend college, but even within this group missingness is substantial. One source of missingness is transcript non-receipt: these variables are derived from the NLSY97 Postsecondary Transcript Study, and are therefore observed only for respondents for whom a transcript was successfully obtained. Transcript non-response is the primary driver of missingness: institutions sometimes did not supply transcripts, supplied incomplete records, or provided degree-program information without course-level grades.101010see NLSY97 Appendix 12: Postsecondary Transcript Study Documentation (https://www.nlsinfo.org/content/cohorts/nlsy97/other-documentation/codebook-supplement/appendix-12-post-secondary-transcript-study). Another potential source of missingness may be that students drop out of college before declaring a major. Consistent with this, among college-goers who do not complete a BA, approximately 10% are missing information on degree major, compared with 1.5% of BA completers. This pattern raises a natural concern for my strategy to impute missing values for STEM. If missingness reflects a student’s failure to declare a major, then this missingness should be reflected as true missingness in an additional level of the STEM variable, rather than being approached as a missing data problem. At the same time, it is difficult to disentangle true non-declaration of a major from missingness caused by institutional non-compliance with transcript requests; students who leave college early may also disproportionately attend institutions that are less responsive to the NLSY’s Postsecondary Transcript Study. As a result, the observed missingness conflates measurement limitations with underlying educational progression in ways that cannot be fully separated.